Prove that Cos^2 x(1-sec^2 x) = -sin^2 x

Answer by mananth(16946)

Cos^2 x(1-sec^2 x) = -sin^2 x

Cos^2 x(Sec^2x-tan^2x-sec^2 x) = -sin^2 x

cos^2x * -tan^2x
cos^2x *-sin^2x/cos^2x
=-sin^2x